OpenStudy (anonymous):

Wendy took a trip from city A to city B, a distance of 390 mi. She traveled part of the way by bus, which arrived at the train station just in time for Wendy to complete her journey by train. The bus averaged 40 mi/h, and the train averaged 85 mi/h. The entire trip took 7.5 hours. How long did Wendy spend on the train?

OpenStudy (anonymous):

t = d/r Solve the following for d, the distance the Wendy was on the train.\[\frac{390-d}{40}+\frac{d}{85}=7.5 \]d = 170\[\frac{170}{85}=2 \text{ hours}\]

OpenStudy (anonymous):

@Freyes Thank you for the medal.

OpenStudy (anonymous):

@Rhope Thanks for the medal.